# Logical perspectives on learning statistical objects

### v1

@@ 6.2 Preservation of online learnability in moving to statistical classes, via logic
(Chase & Freitag, 2019) A formula $\phi(\vec{x},\vec{y})$ is stable if and only if the concept class ${\mathcal{C}}_{\phi}$ is online learnable.
@@
Here we refer to learnability either in the realizable case or the agnostic case, which are equivalent for a concept class, as noted in the preliminaries. Both are equivalent to ${\mathcal{C}}_{\phi}$ having finite Littlestone dimension by Fact 6.
A first-order structure is said to be stable when every partitioned formula is stable. As with NIP, many classical structures are already known to be stable, which implies that all definable families are online learnable. These include the complex field, vector spaces, and commonly studied equivalence relations (Simon, 2015, Example 2.61).
@@
Thus far we are reviewing a connection between the definability property stability and online learnability, which is already known. We now turn to real-valued structures and continuous logic, where to the best of our knowledge, there was no characterization of which structures were online learnable. Recall that in the case of continuous logic structures partitioned real-valued formulas define classes of real-valued functions. The notion of a stable formula generalizes to this setting, but now requires a real parameter:
Fact 24 ((Ben Yaacov & Usvyatsov, 2010, Section 7)).
If $\phi(x;y)$ is a continuous logic formula well-defined over a structure $\mathfrak{C}$ and $\gamma>0$, the following are equivalent:
  • There is some $d$ such that in $\mathfrak{C}$ there are no $a_{1},\dots,a_{d}$, $b_{1},\dots,b_{d}$ such that for all $i<j$,
    $|\phi(a_{i};b_{j})-\phi(a_{j};b_{i})|\geq\gamma$
  • There is some $d$ such that in $\mathfrak{C}$ there are no $a_{1},\dots,a_{d}$, $b_{1},\dots,b_{d}$ and $r,s$ with $r+\gamma\leq s$ such that for all $i<j$, $\phi(a_{i};b_{j})\leq r$ and $\phi(a_{j};b_{i})\geq s$.
@@
We call such a formula $\gamma$-stable. If $\phi(x;y)$ is $\gamma$-stable for all $\gamma>0$, then it is called stable. Roughly speaking stability says that we cannot use gaps in function values, discretized up to some $\gamma$, to define arbitrarily large linear orders.
@@
Theorem 8.
A partitioned formula $\phi(x;y)$ of continuous logic defined over a structure $\mathfrak{C}$ is stable if and only if for every $\gamma>0$, the sequential fat-shattering dimension $\mathrm{FatSHDim}^{Seq}_{\gamma}({\mathcal{H}}_{\phi})$ is finite.
Specifically, if $\phi(x;y)$ is $\delta$-stable for $0<\delta<\frac{\gamma}{2}$, then $\mathrm{FatSHDim}^{Seq}_{\gamma}({\mathcal{H}}_{\phi})$ is finite, and if $\mathrm{FatSHDim}^{Seq}_{\gamma}({\mathcal{H}}_{\phi})$ is finite, then $\phi(x;y)$ is $\gamma$-stable.
The interest in the above theorem is that stability has already been shown to be preserved in moving to from a base class to a statistical class:
Fact 25.
(Ben Yaacov & Keisler, 2009; Ben Yaacov, 2013) If a partitioned CLformula $\phi(\vec{x},\vec{y})$ is stable over a real-valued model $\mathfrak{C}$, then the corresponding formula $E\phi(\vec{X};\vec{Y})$ over the randomization is also stable in the randomization $\mathfrak{C}$.
Thus we can combine Theorem 8, Fact 25, and the characterization of agnostic online learnability via sequential fat-shattering in Fact 6 to give another proof that preservation of agnostic online learning in moving to statistical classes:
Corollary 3.
For any partitioned $\phi(\vec{x};\vec{y})$ over a classical or real-valued model $\mathfrak{C}$, the hypothesis class of $\phi$ is agnostic online learnable if and only the randomized version $E\phi(\vec{X};\vec{Y})$ is agnostic online learnable. In particular, if the hypothesis class associated to the formula is agnostic online learnable, so are the distribution class and the dual distribution class.
Again, we emphasize, that have stated the result here for classes defined by formulas, since this matches the prior development in model theory. But the result extends to general real-valued classes.
@@
Proof.
Assume that $\phi(x;y)$ is not $\gamma$-stable. We will show that ${\mathcal{H}}_{\phi}$ $\gamma$ fat-shatters a binary tree of depth $d$. First, we linearly order the set $\{-1,1\}^{\leq d}$ in a variation of lexicographical fashion, so that for any string $s$ of length $k$ with $k<d$, and any string $t$ strictly extending $s$, if $t_{k}=-1$ then $t<s$, and if $t_{k}=1$ then $t>s$. Thus by instability, there are $(a_{s}b_{s}:s\in\{-1,1\}^{\leq d})$ such that for all $s<t$ in this order, $\phi(a_{s};b_{t})\leq r$ and $\phi(a_{t};b_{s})\geq s.$
We claim we can shatter the tree $T$ defined by sending each sequence $E\in\{-1,1\}^{<d}$ to $a_{E}$. For every $E\in\{-1,1\}^{d}$, let $E_{<t}=(E(0),\dots,E(t-1))$. We consider $b_{E}$, and see that for all $0\leq t<d$, if $E(t)=-1$, then $E<E_{<t}$, so $\phi(a_{E_{<t}};b_{E})\leq r$, while if $E(t)=1$, we have $E>E_{<t}$ and $\phi(a_{E_{<t}};b_{E})\geq s$. As $s-r\geq\gamma$, this tree is $\gamma$-shattered. Thus we have finished the proof of one direction of Theorem 8.
To prove the other direction, we fix $0<\delta<\frac{\gamma}{2}$ and assume that for every $d$, ${\mathcal{H}}_{\phi}$ $\gamma$-shatters a binary tree $T$ in $X$ of depth $d$, and prove that $\phi(x;y)$ is not $\delta$-stable.
To do this, we show by induction on $d$ that if $k>(\gamma-2\delta)^{-1}$ and ${\mathcal{H}}$ is a class on $X$ that $\gamma$-shatters a binary tree $T$ of depth $\frac{k^{d+1}-1}{k-1}$, then there are $x_{1},\dots,x_{d}\in X$ and $h_{1},\dots,h_{d}\in{\mathcal{H}}$ such that for all $i<j$, $|h_{i}(x_{j})-h_{j}(x_{i})|\geq\delta$.
@@
Fact 26 (Alon et al. (2019, Lemma 16)).
@@
Corollary 4 (Tree Ramsey).
@@
Thus we have completed the proof of the other direction of Theorem 8. ∎
@@ 6.3 Realizable Online Learning for statistical classes
We now turn to preservation of realizable online learning for statistical classes.
@@
Fact 27 (Attias et al. (2023, Theorem 4)).
@@
Proof.
Suppose $T$ is the $\gamma$-fat-shattered tree. Then fix a binary tree $s$ of depth $d$ in ${\mathcal{R}}$, and label each branch $b\in\{-1,1\}^{d}$ with some $h_{b}\in{\mathcal{H}}$ such that for all nodes $t$ of the tree $\{-1,1\}^{d}$, $b_{-1},b_{1}$ are branches extending $t$, and $b_{i}$ extends $t$ concatenated with $i$ for $i=\pm 1$, then
@@
where $t_{i}$ is the restriction of $b$ to level $i$. As $\tau$ takes a constant value $\gamma-\epsilon$, the online dimension is greater than $d(\gamma-\epsilon)$. ∎
@@
Because this gives a finite bound for all $\gamma$, realizable online learnability implies agnostic online learnability, even for real-valued function classes.
@@
Proof.
Let $X$ be the infinite binary tree $X=\bigcup_{t=0}^{\infty}\{0,1\}^{t}$. Fix a decreasing sequence $\Gamma=\gamma_{0},\gamma_{1},\dots$ of positive reals to be determined, with $\lim_{d}\gamma_{d}=0$. We define ${\mathcal{H}}^{\Gamma}$, a hypothesis class indexed by the infinite branches $\{0,1\}^{\mathbb{N}}$ of that infinite binary tree.
@@
We now see that by choosing $\Gamma=\gamma_{1},\gamma_{2},\dots$ so that $\lim_{d}d\cdot\gamma_{d}=\infty$ and choosing $\Gamma^{\prime}=\gamma^{\prime}_{1},\gamma^{\prime}_{2},\dots$ so that $\sum_{i=1}^{\infty}\gamma^{\prime}_{i}$ converges, we find ${\mathcal{H}}^{\Gamma}$ with infinite online dimension and ${\mathcal{H}}^{\Gamma^{\prime}}$ with finite online dimension such that $f\circ{\mathcal{H}}^{\Gamma}={\mathcal{H}}^{\Gamma^{\prime}}$. ∎
@@
Proof.
Let $\Gamma=\gamma_{i}:i>0$ be a sequence such that the $i\cdot\gamma_{i}$ is unbounded. Let ${\mathcal{H}}^{\Gamma}$ be the class from Theorem 9. Recall that range points are prefixes $p$ (finite sequences). Hypotheses are parameterized by infinite sequences $s$ and the value of a hypothesis $h_{s}$ on a prefix $p$ is either zero or $\frac{1}{\gamma_{n}}$ for $n$ the length of $p$. Then, as proven in Theorem 9 $H_{\gamma}$ is not online learnable in the realizable case.
@@
To finish the proof, we note that $H_{\gamma}$ is the dual of $D_{\gamma}$. ∎

### v2

@@ 5.2 Preservation of online learnability in moving to statistical classes, via stability
(Chase and Freitag, 2019) A concept class is stable if and only if it is online learnable.
@@
Here we refer to learnability either in the realizable case or the agnostic case, which are equivalent for a concept class, as noted in the preliminaries. Both are equivalent to the class having finite Littlestone dimension by Fact 6.
@@
Thus far we are reviewing a connection between stability and online learnability for concept classes, which is already known. We now turn to real-valued classes. The notion of a stable hypothesis class generalizes to this setting, but now requires a real parameter. It can be defined in terms of either of the following dimensions:
Definition 40
Let ${\mathcal{H}}$ be a hypothesis class on $X$.
For any $\gamma>0$ and any $d$, say that ${\mathcal{H}}$ has $\gamma-$threshold dimension at least $d$ when there are $a_{1},\dots,a_{d}\in X$, $h_{1},\dots,h_{d}\in{\mathcal{H}}$ such that for all $i<j$,
$|h_{j}(a_{i})-h_{i}(a_{j})|\geq\gamma.$
For any $r<s$ and any $d$, say that ${\mathcal{H}}$ has $(r,s)-$threshold dimension at least $d$ when there are $a_{1},\dots,a_{d}\in X$, $h_{1},\dots,h_{d}\in{\mathcal{H}}$ such that for all $i<j$, $h_{j}(a_{i})\leq r$ and $h_{i}(a_{j})\geq s$.
Call a hypothesis class ${\mathcal{H}}$ $\gamma$-stable when it has finite $\gamma$-threshold dimension, and call ${\mathcal{H}}$ $(r,s)$-stable when it has finite $(r,s)$-threshold dimension.
These dimensions give two equivalent definitions of overall stability for a hypothesis class, as shown in the model-theoretic context in (Ben Yaacov and Usvyatsov, 2010, Section 7):
Fact 22 (Ben Yaacov and Usvyatsov (2010, Section 7))
If ${\mathcal{H}}$ is a hypothesis class, the following are equivalent:
  • For every $\gamma>0$, ${\mathcal{H}}$ is $\gamma$-stable.
  • For every $r<s$, ${\mathcal{H}}$ has is $(r,s)-$stable.
@@
We call such a class stable. Roughly speaking stability says that we cannot use gaps in function values, discretized up to some $\gamma$, to define arbitrarily large linear orders.
We will give a slightly more explicit version of this equivalence, using largely the same proof as in Ben Yaacov and Usvyatsov (2010), but with only finitary Ramsey theory, in order to define stability uniformly over a family of hypothesis classes.
Definition 41
If ${\mathcal{F}}$ is a family of hypothesis classes on $X$ parameterized by $Y$, and $\gamma>0$, say that ${\mathcal{F}}$ is uniformly $\gamma$-stable when there is some $d$ such that every ${\mathcal{H}}\in{\mathcal{F}}$ has $\gamma-$threshold dimension at most $d$.
For $r<s$, say that ${\mathcal{F}}$ is uniformly $(r,s)$-stable when there is some $d$ such that every ${\mathcal{H}}\in{\mathcal{F}}$ has $(r,s)-$threshold dimension at most $d$.
These two notions are closely related:
Lemma 42
For any $\gamma>0$ and $d$, if ${\mathcal{H}}$ has $\gamma$-threshold dimension less than $d$, then for all $r,s$ with $r+\gamma\leq s$, ${\mathcal{H}}$ has $(r,s)$-threshold dimension less than $d$.
For any $0<\delta<\gamma$ and $d$, there is $d^{\prime}$ such that if ${\mathcal{H}}$ has $(r,s)$-threshold dimension less than $d$ whenever $r+\delta\leq s$, then ${\mathcal{H}}$ has $\gamma$-threshold dimension less than $d^{\prime}$.
Proof  It is straightforward to see that if ${\mathcal{H}}$ has $(r,s)$-threshold dimension at least $d$, then the same $a_{1},\dots,a_{d}\in X$, $h_{1},\dots,h_{d}\in{\mathcal{H}}$ that recognize this show that ${\mathcal{H}}$ has $(s-r)$-threshold dimension at least $d$.
Now fix $0<\delta<\gamma$ and $d$. Let $n=\lceil\frac{1}{\gamma-\delta}\rceil$. Then by Ramsey’s theorem, there is some $d^{\prime}$ such that any coloring of the complete graph on $d^{\prime}$ vertices with $2n$ colors admits a monochromatic set of size $d$.
We now show that if ${\mathcal{H}}$ has $\gamma$-threshold dimension at least $d^{\prime}$, then there are some $r<s$ with $r+\delta\leq s$ such that ${\mathcal{H}}$ has $(r,s)$-threshold dimension at least $d$.
By our assumption, there are $a_{1},\dots,a_{d^{\prime}}\in X$, $h_{1},\dots,h_{d^{\prime}}\in{\mathcal{H}}$ such that for all $i<j$, $|h_{j}(a_{i})-h_{i}(a_{j})|\geq\gamma$.
Thus for each $i<j$, because $\frac{1}{n}\leq\gamma-\delta$, there is some $0\leq k<n$ such that either $h_{j}(a_{i})\leq\frac{k}{n},-h_{i}(a_{j})\geq\frac{k}{n}+\delta$, or $h_{i}(a_{j})\leq\frac{k}{n},-h_{j}(a_{i})\geq\frac{k}{n}+\delta$. This gives $2n$ possible cases, so we may color the complete graph on $d^{\prime}$ with $2n$ colors such that every edge $(i,j)$ with $i<j$ given a particular color falls into the same case. We may thus find a monochromatic subset $I\subseteq\{1,\dots,d^{\prime}\}$ of size $d$. Let $k$ be the $k$ appearing in the case corresponding to the color of that monochromatic subset, and then let $r=\frac{k}{n},s=\frac{k}{n}+\delta$. Then either for all $i<j$ in $I$, $h_{j}(a_{i})\leq r,-h_{i}(a_{j})\geq s$, in which case we are done, or for all $i<j$ in $I$, $h_{i}(a_{j})\leq r,-h_{j}(a_{i})\geq s$, in which case we only need to reverse the order of our witnesses.
Thus if we universally quantify over all the parameters, these two notions are the same:
Corollary 43
If ${\mathcal{F}}$ is a family of hypothesis classes on $X$ parameterized by $Y$, then the following are equivalent:
  • For every $\gamma>0$, ${\mathcal{F}}$ is uniformly $\gamma$-stable.
  • For every $r<s$, ${\mathcal{F}}$ is uniformly $(r,s)$-stable.
This lets us define two equivalent properties of a class ${\mathcal{F}}$ if either holds, we call ${\mathcal{F}}$ uniformly stable.
@@
Theorem 44
A hypothesis class ${\mathcal{H}}$ is stable if and only if for every $\gamma>0$, the sequential fat-shattering dimension $\mathrm{FatSHDim}^{Seq}_{\gamma}({\mathcal{H}})$ is finite, and a family ${\mathcal{F}}$ of hypothesis classes is uniformly stable if and only if for every $\gamma>0$, there is some $d$ such that for each ${\mathcal{H}}\in{\mathcal{F}},$ the sequential fat-shattering dimension $\mathrm{FatSHDim}^{Seq}_{\gamma}({\mathcal{H}})$ is at most $d$.
Specifically, the following implications hold:
  • If ${\mathcal{H}}$ has $\gamma$ sequential fat-shattering dimension less than $d$, and $r+\gamma\leq s$, then ${\mathcal{H}}$ has $(r,s)$-threshold dimension less than $12^{d+1}-$ 1.
  • If $0<\delta<\frac{\gamma}{2}$, then for every $d$ there is some $d^{\prime}$ such that if ${\mathcal{H}}$ has $\delta$-threshold dimension less than $d$, then ${\mathcal{H}}$ has $\gamma$ sequential fat-shattering dimension less than $d^{\prime}.$
The interest in the above theorem is that stability has already been shown to be preserved in moving to from a measurable family to its expectation class:
Fact 23 (Ben Yaacov and Keisler (2009); Ben Yaacov (2013))
If a measurable family ${\mathcal{F}}$ of hypothesis classes on $X$ parameterized by $Y$ is uniformly stable, then the expectation class ${\mathbb{E}}{\mathcal{F}}$ is also stable.
The result above is phrased very differently in the cited source, in terms of continuous logic. In Appendix D we explain how to get from the statement in the sources to this version.
Thus we can combine Theorem 44, Fact 23, and the characterization of agnostic online learnability via sequential fat-shattering in Fact 6 to give another proof that preservation of agnostic online learning in moving to statistical classes:
Corollary 45
For any measurable family ${\mathcal{F}}$ of hypothesis classes on $X$ parameterized by $Y$, the expectation class ${\mathbb{E}}{\mathcal{F}}$ is agnostic online learnable if and only if for every $\gamma>0$, there is some $d$ such that for each ${\mathcal{H}}$ in the range of ${\mathcal{F}}$, ${\mathcal{H}}$ has $\gamma$ sequential fat-shattering dimension at most $d$.
In particular, if a hypothesis class ${\mathcal{H}}$ is agnostic online learnable, so are its distribution class and dual distribution class.
@@
Proof  Assume that ${\mathcal{H}}$ has $(r,s)$-threshold dimension at least $12^{d+1}-$ 1, with $\gamma=s-r$. We will show that ${\mathcal{H}}$ $\gamma$ fat-shatters a binary tree of depth $d$. First, we linearly order the set $\{-1,1\}^{\leq d}$ in a variation of lexicographical fashion, so that for any string $t_{1}$ of length $k$ with $k<d$, and any string $t_{2}$ strictly extending $t_{1}$, if $(t_{2})_{k}=-1$ then $t_{1}<t_{2}$, and if $(t_{2})_{k}=1$ then $t_{2}>t_{1}$. By the assumption that ${\mathcal{H}}$ has $(r,s)$-threshold dimension at least $12^{d+1}-$ 1, as the set $\{-1,1\}^{\leq d}$ has size $12^{d+1}-$ 1, there are $((a_{t},h_{t}):t\in\{-1,1\}^{\leq d})$ such that for all $i<j$ in this order, $h_{j}(a_{i})\leq r$ and $h_{i}(a_{j})\geq s.$
We claim we can shatter the tree $T$ defined by sending each sequence $E\in\{-1,1\}^{<d}$ to $a_{E}$. For every $E\in\{-1,1\}^{d}$, let $E_{<t}=(E(0),\dots,E(t-1))$. We consider $h_{E}$, and see that for all $0\leq t<d$, if $E(t)=-1$, then $E<E_{<t}$, so $h_{E}(a_{E_{<t}})\leq r$, while if $E(t)=1$, we have $E>E_{<t}$ and $h_{E}(a_{E_{<t}})\geq s$. As $s-r\geq\gamma$, this tree is $\gamma$-shattered. Thus we have finished the proof of the first dimension implication in Theorem 44.
To prove the other direction, we fix $0<\delta<\frac{\gamma}{2}$. We will show by induction on $d$ that if $k>(\gamma-2\delta)^{-1}$ and ${\mathcal{H}}$ $\gamma$-shatters a binary tree $T$ in $X$ of depth $\frac{k^{d+1}-1}{k-1}$, then there are $x_{1},\dots,x_{d}\in X$ and $h_{1},\dots,h_{d}\in{\mathcal{H}}$ such that for all $i<j$, $|h_{i}(x_{j})-h_{j}(x_{i})|\geq\delta$.
@@
Fact 24 (Alon et al. (2019, Lemma 16))
@@
Corollary 47 (Tree Ramsey)
@@
Thus we have completed the proof of the other direction of Theorem 44.
@@ 5.3 Realizable Online Learning for statistical classes
We now turn to preservation of realizable online learning for statistical classes. As with realizable PAC learning, our results will be negative.
@@
Fact 25 (Attias et al. (2023, Theorem 4))
@@
Proof  Suppose $T$ is the $\gamma$-fat-shattered tree. Then fix a binary tree $s$ of depth $d$ in ${\mathcal{R}}$, and label each branch $b\in\{-1,1\}^{d}$ with some $h_{b}\in{\mathcal{H}}$ such that for all nodes $t$ of the tree $\{-1,1\}^{d}$, $b_{-1},b_{1}$ are branches extending $t$, and $b_{i}$ extends $t$ concatenated with $i$ for $i=\pm 1$, then
@@
where $t_{i}$ is the restriction of $b$ to level $i$. As $\tau$ takes a constant value $\gamma-\epsilon$, the online dimension is greater than $d(\gamma-\epsilon)$.
@@
Because this gives a finite bound for all $\gamma$, we see that:
realizable online learnability implies agnostic online learnability, even for real-valued function classes.
@@
Proof  Let $X$ be the infinite binary tree $X=\bigcup_{t=0}^{\infty}\{0,1\}^{t}$. Fix a decreasing sequence $\Gamma=\gamma_{0},\gamma_{1},\dots$ of positive reals to be determined, with $\lim_{d}\gamma_{d}=0$. We define ${\mathcal{H}}^{\Gamma}$, a hypothesis class indexed by the infinite branches $\{0,1\}^{\mathbb{N}}$ of that infinite binary tree.
@@
We now see that by choosing $\Gamma=\gamma_{1},\gamma_{2},\dots$ so that $\lim_{d}d\cdot\gamma_{d}=\infty$ and choosing $\Gamma^{\prime}=\gamma^{\prime}_{1},\gamma^{\prime}_{2},\dots$ so that $\sum_{i=1}^{\infty}\gamma^{\prime}_{i}$ converges, we find ${\mathcal{H}}^{\Gamma}$ with infinite online dimension and ${\mathcal{H}}^{\Gamma^{\prime}}$ with finite online dimension such that $f\circ{\mathcal{H}}^{\Gamma}={\mathcal{H}}^{\Gamma^{\prime}}$.
@@
Proof  Let $\Gamma=\gamma_{i}:i>0$ be a sequence such that the $i\cdot\gamma_{i}$ is unbounded. Let ${\mathcal{H}}^{\Gamma}$ be the class from Theorem 52. Recall that range points are prefixes $p$ (finite sequences). Hypotheses are parameterized by infinite sequences $s$ and the value of a hypothesis $h_{s}$ on a prefix $p$ is either zero or $\frac{1}{\gamma_{n}}$ for $n$ the length of $p$. Then, as proven in Theorem 52 $H_{\gamma}$ is not online learnable in the realizable case.
@@
To finish the proof, we note that $H_{\gamma}$ is the dual of $D_{\gamma}$.